# Linear Magnetoresistance and Anomalous Hall Effect in the Superconductor NiBi3

### v1

@@ Acknowledgements.
- Hayashi and Ando [2024] H. Hayashi and K. Ando, Two-dimensional rashba superconductivity in ni/bi bilayers evidenced by nonreciprocal transport, Applied Physics Reviews 11, 011401 (2024), https://pubs.aip.org/aip/apr/article-pdf/doi/10.1063/5.0158237/18283582/011401_1_5.0158237.pdf .
@@
First derivative of resistivity is shown in Fig. SM1 for S1 and S3. The divergence in the blue-shade region correspond to the superconducting transition. Above the critical temperature, both samples follows the standard Bloch-Gruneisen (BG) behavior, as indicated by the solid colored lines. The crossover in which phonon modes become thermally excited are represented by the peak in d/dT. For S3, it manifest around 15 K, similar to S2 in the main text, while in S3, the peak appears at 30 K. After the maxima, decreasing of d/dT reflect a deviation from BG predictions of linear-scatter contribution. Such a deviation occurs smoothly for S3 and abruptly for S1, similar to S2. This difference is associated with variation of th crystal size and shape, which is reflect by the difference in Debye temperature (), being around 70 K for S1 and S2, and larger for S3 (85 K).
@@
Magnetization measurements were performed on a batch of NiBi3 single crystals using a Superconducting Quantum Interference Device (SQUID) magnetometer from Quantum Design. Fig. SM2a shows the DC magnetic susceptibility as a function of temperature under various applied magnetic fields. The upper branches correspond to field-cooled (FC) measurements, while the lower branches represent zero-field-cooled (ZFC) data. A noticeable bifurcation between FC and ZFC curves marks the onset of magnetic irreversibility, with the irreversibility temperature for each field indicated by dashed lines. Small oscillations with amplitudes on the order of 1×10-9 emu are attributed to minor variations in sample positioning during measurement and do not reflect intrinsic properties of the crystals. Fig. SM2b presents the magnetization as a function of temperature under an applied field of 500 Oe, clearly demonstrating the diamagnetic response of the sample.
Magnetoresistance (MR) measurements as a function of magnetic field at various temperatures for sample S3 are shown in Fig. SM3a. The out-of-scale plots emphasize the evolution of the MR behavior. From 5 K to 15 K, the MR exhibits a purely quadratic dependence, consistent with ordinary MR mechanisms. At 20 K, a weak linear contribution begins to emerge and becomes increasingly prominent at higher temperatures. As shown in Fig. SM3b, where the MR at 2 T is plotted versus temperature, this linear component reaches a maximum around 75 K and subsequently diminishes. Consequently, as depicted in Fig. SM3c, the overall MR increases with temperature despite a suppression of the quadratic contribution, due to the dominance of the linear term at intermediate temperatures

### v2

@@ Acknowledgements.
- Hayashi and Ando [2024] H. Hayashi and K. Ando, Two-dimensional rashba superconductivity in ni/bi bilayers evidenced by nonreciprocal transport, Applied Physics Reviews 11, 011401 (2024).
@@
- Feng et al. [2019] Y. Feng, Y. Wang, D. M. Silevitch, J.-Q. Yan, R. Kobayashi, M. Hedo, T. Nakama, Y. Ōnuki, A. V. Suslov, B. Mihaila, P. B. Littlewood, and T. F. Rosenbaum, Linear magnetoresistance in the low-field limit in density-wave materials, Proceedings of the National Academy of Sciences 116, 11201 (2019), https://www.pnas.org/doi/pdf/10.1073/pnas.1820092116 .
- Abe et al. [2024] N. Abe, Y. Hano, H. Ishizuka, Y. Kozuka, T. Tadano, Y. Tsujimoto, K. Yamaura, S. Ishiwata, and J. Fujioka, Large anomalous hall effect in spin fluctuating devil’s staircase, npj Quantum Materials 9 (2024).
- Tang et al. [2024] K. Tang, Y. Yang, J. Shen, M. Shi, N. Zhang, H. Li, H. Li, L. Zhengtai, D. Shen, R. Wang, Y. Gao, J. He, Z. Xiang, and X. Chen, Unconventional anomalous hall effect and large anomalous nernst effect in antiferromagnet smmnbi2, Communications Materials 5 (2024).
@@
The carrier density for the NiBi3 is presented in Fig. SM2. The carrier density was obtained by linear fitting the Rxy vs. B curves at high field, where the AHE are saturated. The curve displays a fast decay of carrier density, reaching to a minima at about 50 K. This low temperature carrier suppression can be related to spin fluctuations in the system.
Magnetization measurements were performed on a bunch of NiBi3 single crystals using a Superconducting Quantum Interference Device (SQUID) magnetometer MPMS-XL from Quantum Design. Fig. SM3a shows the DC magnetic susceptibility as a function of temperature under various applied magnetic fields. The upper branches correspond to field-cooled (FC) measurements, while the lower branches represent zero-field-cooled (ZFC) data. A noticeable bifurcation between FC and ZFC curves marks the onset of magnetic irreversibility, with the irreversibility temperature for each field indicated by dashed lines. Small oscillations with amplitudes on the order of 1×10-9 emu are attributed to minor variations in sample positioning during measurement and do not reflect intrinsic properties of the crystals. Fig. SM3b presents the magnetization as a function of temperature under an applied field of 500 Oe, clearly demonstrating the diamagnetic response of the sample.
Magnetoresistance (MR) measurements as a function of magnetic field at various temperatures for sample S3 are shown in Fig. SM4a. From 5 K to 15 K, the MR exhibits a purely quadratic dependence, consistent with ordinary MR mechanisms. At 20 K, a weak linear contribution begins to emerge and becomes increasingly prominent at higher temperatures. As shown in Fig. SM4b, where the MR at 2 T is plotted versus temperature, this linear component reaches a maximum around 75 K and subsequently diminishes. Consequently, as depicted in Fig. SM4c, the overall MR increases with temperature despite a suppression of the quadratic contribution, due to the dominance of the linear term at intermediate temperatures.
Magnetoresistance for Two-Bands Parallel Conduction with Spin Mixing
Following the seminal work of Fert and Campbell on spin-dependent transport in ferromagnetic alloys [43], the total electrical resistivity in the presence of spin-mixing processes can be expressed as
where and denote the partial resistivities associated with spin-up and spin-down conduction channels, and accounts for spin-flip scattering events that couple the two spin channels.
We define the resistivity spin-channel asymmetry ratio
Suppose that Spin mixing resistivity is proportional to the applied magnetic field (H), = H. Therefore the magnetoresistance can be calculated as
After some algebra, one can obtain
Rewriting in terms of , we can simplify to:
where
For low magnetic fields (), Eq. 5 can be expanded, leading to
Considering only the linear term, the magnetoresistance becomes
The central hypothesis in this modified Campbell-Fert model is the assumption of a linear relation between the mixing resistivity term and the external magnetic field. Qualitatively, we argue that in the fluctuation regime, spin fluctuation arises from a non-magnetic ground state, on a time scale larger than the carriers scattering. Magnetic correlation between the fluctuations separates the conduction channels into two components, giving rise to a short-range magnetic ordering which spam the AHE and LMR. The external field promotes an asymmetrical enhancement of the spin-flip by forcing the fluctuating moments to align with the field and increasing the mixing resistivity term.
30 K has = 5.04 and = 131.3 150 K has and